# Space Group Informed Transformer for Crystalline Materials Generation

### v1

@@ Appendix B Stable crystal samples
To estimate the energy above hull, the DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [96] and all-electron projector-augmented wave method [97], as implemented in the VASP code [98]. All parameters of the calculations including settings of PBE functional, Hubbard U corrections, and ferromagnetic initialization are chosen to be consistent with Materials Project by using of MPRelaxSet function in pymatgen [60] A double relaxation strategy was employed. The maximum optimization ionic step and the maximum running time were constrained to 150 steps and 20 hours, respectively. All structures containing Yb element are ignored when calculating energy above hull due to they are unavailable from the Materials Project at the time of writing.
| In MP | Not in MP | ||||
@@
| Formula | (eV/atom) | Lattice constant () | Formula | (eV/atom) | Lattice constant () |
| PdPtTm | -0.00164 | 6.90 (6.86) | AcCuRh | -0.14248 | 7.47 |
| LiPtZr | -0.01668 | 6.43 (6.39) | AcAgSi | -0.04364 | 7.81 |
| CdLiPb | -0.00313 | 6.82 (6.79) | IrLuPm | -0.0289 | 6.89 |
| LiNdPb | -0.05703 | 7.05 (7.03) | InPmTl | -0.01318 | 7.73 |
| BeOsSi | -0.00779 | 5.77 (5.74) | PdPmZn | -0.0843 | 7.27 |
| CeAs | -0.0073 | 5.99 (6.09) | AuInPm | -0.0568 | 7.70 |
| Xe | -0.00067 | 7.45 (6.66) | AcBiK | -0.20215 | 8.37 |
| InLiY | -0.17504 | 7.54 | InTb | -0.26278 | 7.95 |
| InNd | -0.21393 | 8.35 | AcHgIn | -0.40448 | 8.66 |
| BeC | -0.01227 | 4.33 (4.32) | CaPaRu | -0.09344 | 6.80 |
| EuTl | -0.1054 | 7.90 | |||

### v2

@@ Appendix C Discovered crystal samples with symmetric structure initialization
To relax samples and estimate the energy above the hull, the DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [125] and all-electron projector-augmented wave method [126], as implemented in the VASP code [127]. All parameters of the calculations including settings of PBE functional, Hubbard U corrections, and ferromagnetic initialization are chosen to be consistent with Materials Project by using of MPRelaxSet function in pymatgen [81] A double relaxation strategy was employed. The maximum optimization ionic step and the maximum running time were constrained to 150 steps and 20 hours, respectively. All structures containing Yb element are ignored when calculating energy above the hull due to they are unavailable from the Materials Project at the time of writing.
| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) |
| LuSiNi | 62 | c-Lu-c-Si-c-Ni | 0.0134 |
| Eu3Mg | 62 | c-Eu-c-Mg-d-Eu | 0.0661 |
| EuAgAu | 62 | c-Eu-c-Ag-c-Au | 0.0111 |
| SmSbPd | 62 | c-Sm-c-Sb-c-Pd | 0.0368 |
| Nd3Rh | 62 | c-Nd-c-Rh-d-Nd | 0.0673 |
| SrSnPd | 62 | c-Sr-c-Sn-c-Pd | -0.0531 |
| SrCdSn | 62 | c-Sr-c-Cd-c-Sn | 0.0863 |
| SrMgHg | 62 | c-Sr-c-Mg-c-Hg | 0.0517 |
| Eu2Ru | 62 | c-Eu-c-Eu-c-Ru | 0.0890 |
| PrHgAu2 | 62 | c-Pr-c-Hg-c-Au-c-Au | 0.0868 |
| Ce(SiPt)2 | 139 | a-Ce-d-Si-e-Pt | 0.0648 |
| Dy(AlGa)2 | 139 | a-Dy-d-Al-e-Ga | 0.0902 |
| Nd3Dy | 139 | a-Dy-b-Nd-d-Nd | 0.0690 |
| Pd3Pt | 139 | a-Pt-b-Pd-d-Pd | -0.0298 |
| LaMgSn | 139 | c-Mg-e-La-e-Sn | 0.0274 |
| Rb3Mn2F7 | 139 | a-F-b-Rb-e-Rb-e-Mn-e-F-g-F | 0.0874 |
| AlAgAu2 | 139 | a-Al-b-Ag-d-Au | 0.0912 |
| Ba3Tl | 139 | a-Tl-b-Ba-d-Ba | 0.0933 |
| CsPrTe2 | 166 | a-Cs-b-Pr-c-Te | 0.0977 |
| Sr2Ca | 166 | a-Ca-c-Sr | 0.0771 |
| Pr2AlRu3 | 166 | a-Al-c-Pr-d-Ru | 0.0909 |
| Dy3Ho | 194 | d-Ho-h-Dy | 0.0738 |
| Zr3Pb | 194 | d-Pb-h-Zr | 0.0426 |
| PrAu3 | 194 | d-Pr-h-Au | 0.0696 |
| Pd3Pt | 194 | d-Pt-h-Pd | 0.0053 |
| TbZnGa | 194 | a-Tb-c-Ga-d-Zn | 0.0718 |
| BaSr(GaGe)2 | 194 | a-Ba-b-Sr-f-Ga-f-Ge | 0.0855 |
| Ho2Er | 194 | b-Er-f-Ho | 0.0790 |
| NdDyHg2 | 225 | a-Nd-b-Dy-c-Hg | 0.0974 |
| MgAgPd2 | 225 | a-Mg-b-Ag-c-Pd | 0.0540 |
| PrEuIn2 | 225 | a-Eu-b-Pr-c-In | 0.0957 |
| TaNbRu2 | 225 | a-Ta-b-Nb-c-Ru | 0.0823 |
| PdAu | 221 | a-Au-b-Pd | 0.0025 |
| Gd2HgAu | 225 | a-Hg-b-Au-c-Gd | 0.0522 |
1
The fractional coordinates and lattice parameters are omitted for brevity. See details at https://drive.google.com/file/d/1gOIkWkjSH_Ed0-wzPk8VgxLjJcidbrH3/view?usp=sharing
| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) |
| Ba2Si | 62 | b-Ba-c-Ba-c-Si | 0.0766 |
| GdIr | 62 | c-Gd-c-Ir | -0.2310 |
| Nd3Dy | 139 | a-Dy-b-Nd-d-Nd | 0.0680 |
| Pd3Pt | 139 | a-Pd-b-Pt-d-Pd | 0.0147 |
| PdPt | 221 | a-Pd-b-Pt | 0.0847 |
| NdDyHg2 | 225 | a-Dy-b-Nd-c-Hg | 0.0973 |
| MgAgPd2 | 225 | a-Mg-b-Ag-c-Pd | 0.0480 |
| HfRe | 221 | a-Re-b-Hf | 0.0926 |
| TaNbRu2 | 225 | a-Nb-b-Ta-c-Ru | 0.0770 |
| PdAu | 221 | a-Au-b-Pd | 0.0056 |
| Gd2HgAu | 225 | a-Hg-b-Au-c-Gd | 0.0476 |
| TbMgPd2 | 225 | a-Mg-b-Tb-c-Pd | 0.0676 |
1
The fractional coordinates and lattice parameters are omitted for brevity. See details at https://drive.google.com/file/d/18sdsp-6yRSBaNez1A0lr20Ru6H7OOzvJ/view?usp=drive_link.
Appendix D Discovered crystal samples with element substitution
Tables LABEL:table:mutation_samples and LABEL:table:mutation_pmg_samples list double perovskites crystals in the (No. 225) space group with eV/atom found by CrystalFormer and element substitution of pymatgen respectively. We found that CrystalFormer has found three times more perovskites crystals compared to ionic substitution [46]. The DFT calculation setup is the same as in the previous section.
@@ Appendix D Discovered crystal samples with element substitution
| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) | Lattice constant () |
| Sr2TaNiO6 | 225 | a-Ni-b-Ta-c-Sr-e-O | 0.0426 | 7.9762 |
| Ba2TaNiO6 | 225 | a-Ni-b-Ta-c-Ba-e-O | 0.0711 | 8.2129 |
| Ba2YMnO6 | 225 | a-Y-b-Mn-c-Ba-e-O | 0.0284 | 8.3575 |
| Eu2CoMoO6 | 225 | a-Co-b-Mo-c-Eu-e-O | 0.0720 | 7.9220 |
| Ba2TbFeO6 | 225 | a-Tb-b-Fe-c-Ba-e-O | 0.0343 | 8.3200 |
| Ba2HfWO6 | 225 | a-W-b-Hf-c-Ba-e-O | 0.0939 | 8.3568 |
| Eu2CoNiO6 | 225 | a-Ni-b-Co-c-Eu-e-O | 0.0646 | 7.6181 |
| Sr2NbVO6 | 225 | a-V-b-Nb-c-Sr-e-O | 0.0166 | 8.0375 |
| Sr2NaWO6 | 225 | a-Na-b-W-c-Sr-e-O | 0.0901 | 8.2959 |
1
The fractional coordinates are omitted for brevity.
| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) | Lattice constant () |
| Ba2FeBiO6 | 225 | a-Bi-b-Fe-c-Ba-e-O | -0.0030 | 8.3808 |
| Ba2NpCoO6 | 225 | a-Np-b-Co-c-Ba-e-O | 0.0241 | 8.3983 |
| Ba2BiMoO6 | 225 | a-Bi-b-Mo-c-Ba-e-O | 0.0706 | 8.6322 |
1
The fractional coordinates are omitted for brevity.
Appendix E Details of plug-and-play materials design
We build two independent predictive models of band gap and formation energy following the generative pretraining strategy [128]. Generative pretrain of CrystalFormer allows the model to extract crystal representations. Then, by attaching a regression neural network to its feature and training on property labels, the model can be used to predict the properties of crystals.
For the regression task, the mean absolute error corresponds to a predictive model characterized by the Laplace distribution , where is a positive scale parameter. According to Bayes’ rule, the posterior probability for crystals given the property is:
The crystal log-likelihood with a given property condition is then given by:
It is instructive to see that property-guided samples are balanced between the CrystalFormer prior and the property prediction model likelihood. Here plays the role of guidance strength. For multi-property prediction, the likelihood function can be extended to:
Note that MCMC sampling via the Metropolis algorithm does not involve the normalization factor of the conditional distribution. In our experiments, we set for band gap prediction and for formation energy prediction, corresponding to the inverses of the MAE values obtained from the regression model.